OpenStudy (anonymous):

x/2-y/3=5/6 x/5-y/4=-13/10

OpenStudy (campbell_st):

are there 2 equations here? \[\frac{x}{2} - \frac{y}{3} = \frac{5}{6}\] and \[\frac{x}{5} - \frac{y}{4} = \frac{-13}{10}\] and what do you need to do to the equations..?

OpenStudy (anonymous):

substitute